# Integrable Cosmological Models with an Arbitrary Number of Scalar Fields

### v1

@@ Abstract
We consider cosmological models with an arbitrary number of scalar fields nonminimally coupled to gravity and construct new integrable cosmological models. In the constructed models, the Ricci scalar is an integral of motion irrespectively of the type of metric. The general solutions of evolution equations in the spatially flat FLRW metric have been found for models the quartic potentials.
@@ 1 Introduction
In FLRW metric, the Einstein’s equations reduce to a system of ordinary differential equations, but this system is integrable only for a few specific forms of the potential. The problem of constructing integrable cosmological models is actively investigated, but even taking into account one-field cosmological models, the number of integrable models is not large [1, 2, 3, 4, 5, 6, 7, 8, 9]. In the case of a few scalar fields, even construction of cosmological models with exact particular solutions is not a trivial task [10, 11, 12, 13, 14, 15, 16, 17, 18, 19].
There are a few methods for integrating equations of evolution. The integrability of many cosmological models has been proven by solving evolution equations with suitable parametric time [4], but it is not clear how to find this parametric time. A method for constructing integrable models with non-minimally coupled scalar fields by using the relation between the Jordan and Einstein frames has been investigated for one-field models in Refs. [5, 7, 20]. The integrability of cosmological models with non-minimal coupling can be more apparent than the integrability of the corresponding models in the Einstein frame [7].
@@ 2 Model with $N$ scalar fields and nonminimal coupling
Let us consider the model described by the following action
@@ 3 General solutions in the FLRW metric
    $H(t)=\sqrt{\frac{R_{0}}{12}}\coth\left(\sqrt{\frac{R_{0}}{3}}(t-t_{0})\right),$ (25)
    $a(t)=a_{0}\sqrt{\sinh\left(\sqrt{\frac{R_{0}}{3}}(t-t_{0})\right)},$ (26)
@@
    $H(t)=\sqrt{\frac{R_{0}}{12}}\tanh\left(\sqrt{\frac{R_{0}}{3}}(t-t_{0})\right),$ (29)
    $a(t)=a_{0}\sqrt{\cosh\left(\sqrt{\frac{R_{0}}{3}}(t-t_{0})\right)},$ (30)
@@
    $H(t)=\frac{1}{2(t-t_{0})},\qquad a(t)=a_{0}\sqrt{t-t_{0}},$ (35)
@@
$K(m)\equiv F(\pi/2\left|m\right.)$ is the complete elliptic integral of the first kind, and also the Jacobi elliptic functions $\mathrm{sn}$, $\mathrm{cn}$ and $\mathrm{dn}$ were introduced. For reference, here are the definitions of these functions: if $u=F\left(\left.\varphi\,\right|m\right)$, then
@@
    Here $u_{A}=|c_{A}/C_{A}|^{1/4}y^{A}(\eta_{i})$, $\eta_{i}$ is a constant, and the choice of sign in “$\pm$” is determined by initial conditions. Note that there is no summation by $A$.
@@
    There are no solutions when $C^{A}<0$ and $c_{A}\geq 0$.
@@ 4 Conclusions
New integrable models with minimally coupled scalar fields can be obtained by conformal metric transformation of integrable models with non-minimally coupled scalar fields [25]. Another possible way to generalize the obtained results is to obtain the general solutions of integrable FLRW models in open and closed FLRW metrics [7, 26] or in the Bianchi I metric [20, 8, 27].

### v2

@@ Abstract
We consider cosmological models with an arbitrary number of scalar fields nonminimally coupled to gravity and construct new integrable cosmological models. In the constructed models, the Ricci scalar is an integral of motion irrespectively of the type of metric. The general solutions of evolution equations in the spatially flat FLRW metric have been found for models with the quartic potentials.
@@ 1 Introduction
In FLRW metric, the Einstein’s equations reduce to a system of ordinary differential equations, but this system is integrable only for a few specific forms of the potential. The problem of constructing integrable cosmological models is actively investigated, but even taking into account single-field cosmological models, the number of integrable models is not large [1, 2, 3, 4, 5, 6, 7, 8, 9]. In the case of a few scalar fields, even construction of cosmological models with exact particular solutions is not a trivial task [10, 11, 12, 13, 14, 15, 16, 17, 18, 19].
There are a few methods for integrating equations of evolution. The integrability of many cosmological models has been proven by solving evolution equations with suitable parametric time [4], but it is not clear how to find this parametric time. A method for constructing integrable models with non-minimally coupled scalar fields by using the relation between the Jordan and Einstein frames has been investigated for single-field models in Refs. [5, 7, 20]. The integrability of cosmological models with non-minimal coupling can be more apparent than the integrability of the corresponding models in the Einstein frame [7].
@@ 2 Model with $N$ scalar fields and nonminimal coupling
Let us consider the model, described by the following action
@@ 3 General solutions in the FLRW metric
    $\begin{split}&H(t)=\sqrt{\frac{R_{0}}{12}}\coth\left(\sqrt{\frac{R_{0}}{3}}(t-t_{0})\right),\\ &a(t)=a_{0}\sqrt{\sinh\left(\sqrt{\frac{R_{0}}{3}}(t-t_{0})\right)},\end{split}$ (25)
@@
    $\begin{split}&H(t)=\sqrt{\frac{R_{0}}{12}}\tanh\left(\sqrt{\frac{R_{0}}{3}}(t-t_{0})\right),\\ &a(t)=a_{0}\sqrt{\cosh\left(\sqrt{\frac{R_{0}}{3}}(t-t_{0})\right)},\end{split}$ (28)
@@
    $H(t)=\frac{1}{2(t-t_{0})},\quad a(t)=a_{0}\sqrt{t-t_{0}},$ (33)
@@
$K(m)\equiv F(\pi/2\left|m\right.)$ is the complete elliptic integral of the first kind, and also the Jacobi elliptic functions $\mathrm{sn}$, $\mathrm{cn}$ and $\mathrm{dn}$ were introduced. By definition, if $u=F\left(\left.\varphi\,\right|m\right)$, then
@@
    Constants $u_{A}=|c_{A}/C_{A}|^{1/4}y^{A}(\eta_{i})$ and the choice of sign in “$\pm$” are determined by initial conditions set at $\eta=\eta_{i}$. Note that no summation over $A$ is assumed.
@@
    There are no solutions when $C^{A}<0$ and $c_{A}\geqslant 0$.
@@ 4 Conclusions
New integrable models with minimally coupled scalar fields can be obtained by conformal metric transformation of integrable models with non-minimally coupled scalar fields [25]. Another possible way to generalize the obtained results is to get the general solutions of integrable FLRW models in open and closed FLRW metrics [7, 26] or in the Bianchi I metric [8, 20, 27].